there are 1200 soldiers in a fort and there has enough food for 28 day. after 4 day some soldiers transferred hance, food lasted for 32 days. how many soldiers were transferred?
Answer by ankor@dixie-net.com(22740)

there are 1200 soldiers in a fort and there has enough food for 28 days.
after 4 days some soldiers transferred hence, food lasted for 32 days.
how many soldiers were transferred?
:
Find how many food-days available: 1200 *28 = 33,600 food-days
:
let m = no. of men transferred
4days at full strength + 28 days at reduced strength
1200*4 + 28(1200-m) = 33600
4800 + 33600 - 28m = 33600
38400 - 28m = 33600
38400 - 33600 = 28m
4800 = 28m
m = 4800/28
m = 171 men. were transferred